A store buys a product from a supplier for $572 per unit and sells it in the store. The store has a yearly fixed cost of $63615. The product demand is estimated to be 958 units per year. The store has an income tax rate of 31%. Find the unit selling price if the store would like to have a yearly after-tax income of $54837.
Answer #1

Pre-tax income => After tax income/(1-tax rate) = 54837/(1-31%) = 79473.91304

To achieve Pre-tax income 79473.91304 let, the unit selling price required = x

So, Revenue-(fixed cost+variable cost) = pre-tax income

or, 958*x-63615-958*572 = 79473.91304

or, x = (79473.91304+63615+958*572)/958 = 721.3621222 (Ans)